# Topological Structure Learning Should Be A Research Priority for LLM-Based Multi-Agent Systems

### v1

@@ 3.2 Structure Profiling
Structure profiling seeks to determine the macro-level pattern of how agents should be connected to effectively coordinate and solve tasks. However, fixed structure, such as chain, star, or fully connected graph, often fails to generalize across diverse tasks. The optimal macro-level structure pattern is typically task-specific and implicit, depending on factors like reasoning complexity, agent roles, and task demands. For instance, sequential workflows (e.g., multi-step reasoning) align well with chains, while decentralized exploration may favor tree structures. One pressing open problem is to identify the optimal macro-level communication pattern among agents in MASs that maximizes task performance while minimizing unnecessary communication overhead.
Position. We argue that adaptive topology policies that generate a bespoke graph for each query $x$ are now essential.
This section studies how selected agents are connected, communicate, and divide tasks with each other to maximize task reward while controlling communication cost. Formally, a structure is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ over the agent pool $\mathcal{A}$, where an edge $e_{ij}\in\mathcal{E}$ indicates agent $a_{i}$’s output is forwarded as input context to agent $a_{j}$, namely $a_{i}\to a_{j}$. Designing an optimal agent interaction structure is non-trivial. Manually crafting a fixed communication or workflow pattern, such as a chain-of-thought style relay between agents, may work for some tasks but fails to generalize across diverse problems [23]. The appropriate structure is often task-dependent, e.g., a simple question might only require a single agent, whereas a complex query might benefit from a divide-and-conquer approach with multiple specialists. This dependency creates a need for multi-agent structure profiling, i.e., determining the best interaction architecture for a given task.
We define our objective by the utility function $\mathcal{U}$ defined in Section 3.1, which measures the quality of result (and possibly efficiency) when using structure $\mathcal{G}$ on task $x$. This could be, for instance, the task reward or accuracy achieved by the agents minus a penalty for communication cost (number of messages or tokens exchanged). The goal of structure profiling is to find the optimal agent network $\mathcal{G}$ for each task. Formally, for a given query $x$ we seek:
$\mathcal{G}^{*}(x)\;=\;\arg\max_{\mathcal{G}\in\mathcal{S}}\;\mathcal{U}(\mathcal{G};x).$ (6)
@@
Automatically profiling and selecting an effective structure for agents rather than relying on ad-hoc human design is thus a critical open problem.
Pre-defined structures give useful inductive biases but lack flexibility. Meta-agent planners such as AOP [24] and ReMA [25] add task awareness, yet still rely on manually crafted rule sets. Graph-generation approaches, such as G-Designer [26] and CommFormer [27]), demonstrate that fully learned topologies can beat fixed ones, but they target narrow domains or small agent pools. We will propose SPAN, a task-conditioned and scalable network.
@@
Model the macro-structure determination problem as a probabilistic graph.
Existing work shows that the same agents succeed or fail relying on their connection structure. A fully-connected graph outperforms a chain on open-ended research questions [28], whereas a lightweight DAG scales better than regular patterns when hundreds of agents cooperate [12]. Fixed, hand-designed patterns break down as tasks diversify [23]. Therefore, we propose our policy SPAN (Structure Profiling Adaptive Network). The SPAN involves a task encoder, an agent-subset selector, and an edge probability predictor. An encoder will map the raw task description, e.g., the user prompt plus optional metadata to a continuous latent vector. It may be instantiated as a frozen foundation model or a shallow adapter network on top of an LLM prompt embedding. Its sole role is to provide a compact task representation that conditions every subsequent decision. Given task representation and possible learned embeddings from Section 3.1 for candidate agents, the selector may output the Bernoulli probabilities to filter an agent subset by a deterministic threshold. Finally, the key idea is a probabilistic adjacency matrix for edge probability prediction:
$P(x)\in[0,1]^{N\times N},\quad P_{ij}(x)=\Pr(a_{i}\!\to a_{j}\mid x).$ (7)
Given a new task, the policy scores every possible communication link. Thresholding or sampling $P(x)$ instantiates a discrete directed graph, which may be sparse or irregular, always tailored to the task. The probabilities in $P(x)$ factorize decisions, keeping search linear in the number of potential edges rather than exponential in whole graphs. This abstraction (1) unifies prior work that searches end-to-end workflows [29], evolves graphs [23], or routes queries dynamically [16]; and (2) exposes continuous parameters that can be optimized by gradient or RL methods. More SPAN details and related work in Appendix D.
Open problem 4.
How can a society of agent self-edit its topology while a dialogue is still in flight, without creating instability or runaway costs?
While we can select a suitable macro-structure, the agents involved in is actually dynamically changing, i.e., the macro-structure could also be changing due to newly added or deleted agent nodes. In this case, there should be new research problems for macro-structure adaptation.
@@
It is possible to design local edge policies that treat each message as evidence for edge build or break, and meta agents monitor utility gradients and trigger Markov Chain Monte Carlo (MCMC) moves, i.e., add, drop, swap, subject to acyclicity constraints.
In addition, it is hard to compare structural performance and innovations without a public benchmark. So, it is valuable to curate an open suite that pairs each task with a groundtruth or at least oracle-verified near-optimal topology, plus cost models for token, latency, and energy usage, as well as provide diagnostic leader boards that break out performance by macro-structure class.
@@ 3.3 Topological Structure Synthesis
While agent selection and structure profiling identify the macro-level pattern of interactions, synthesizing a micro-level communication topology is a pressing problem to further enhance the communication efficiency and task performance. Specifically, agent selection determines who participates in the collaboration, and structure profiling identifies the macro-level pattern of interaction. Effectively executing the task requires synthesizing a fine-grained communication topology that governs the actual flow of information between agents. This micro-level structure plays a critical role in determining not only which agents should communicate but also the edge directions, weights, and potential hierarchy (e.g., tree depth or breadth). These decisions directly influence how efficiently information propagates throughout the system, ultimately shaping task performance [14], communication cost [14, 13], and the robustness of the multi-agent collaboration [15].
Open problem 5.
Fully connecting every pair of agents leads to redundancy and high communication costs: How do we prune unnecessary links while preserving task performance?
Recent efforts have explored topological structure optimization in MASs through topology pruning-based methods, which aim to reduce redundancy and improve efficiency by selectively removing communication links from the topological structure. AgentPrune [13] learns a sparse spatial-temporal communication graph by optimizing differentiable edge masks with policy gradients and performs one-shot pruning to remove redundant message-passing edges while preserving overall task performance. Similarly, G-Designer [26] adopts a reinforcement learning-based framework to learn task-specific sparse communication graphs by selecting critical communication links. AgentDropout [14], in contrast, adaptively drops low-utility intra- and inter-round edges to compress communication graphs and reduce token usage during inference. GPTSwarm [19] regards agents as nodes in a computational graph and optimizes inter-agent communication links, enabling end-to-end learning of sparse collaboration topologies tailored to task requirements.
@@
Different agent roles should occupy specific positions within the topological structure: How do we design the micro-structure while considering the specific role of each agent to maximize performance?
While many existing approaches focus on the communication connectivity among agents, they often ignore the problem of the different, specialized roles of agents that should occupy distinct and meaningful positions within the topology. For example, a manager should be in the central position for overall coordination, while a verifier may serve best at the end of the pipeline.
Position. We advocate that we should further optimize the macro-level topological patterns in Sec. 3.2 and provide a micro-level topological structure in MASs.
Formulation. The objective of topology synthesis is to synthesize an optimized directed graph $\mathcal{G}=(\hat{\mathcal{A}},\mathcal{E})$ that concretely realizes the macro-structure $M$ over the selected agents $\hat{\mathcal{A}}$. Let $\mathcal{G}_{M}(\hat{\mathcal{A}})$ denote the space of valid topologies consistent with macro-structure $M$. For each candidate graph $\mathcal{G}\in\mathcal{G}_{M}(\hat{\mathcal{A}})$, we define $\mathcal{R}(\mathcal{T},\mathcal{G})$ as the expected task reward and $C(G)$ as the total communication cost. We then solve:
$\mathcal{G}^{*}=\arg\max_{\mathcal{G}\in\mathcal{G}_{M}(\hat{\mathcal{A}})}\mathcal{R}(\mathcal{T},\mathcal{G})-\lambda\cdot C(\mathcal{G}).$ (8)
@@
A central challenge in optimizing micro-level topologies within MASs lies in understanding the specific contribution of each communication link to task performance. Unlike coarse-grained structural assumptions, such as fully connected or fixed topologies, fine-grained assessment allows the system to tailor communication pathways to the demands of each task. One promising approach is counterfactual reasoning, where the utility of a given link is estimated by observing the performance drop caused by its removal. Identifying the essential links or redundant communication links forms pruning strategies that retain only high-utility connections, thereby reducing communication overhead without compromising performance. To further mitigate the high computational cost of evaluating the effects of each communication link, approximation techniques need to be used for scalability. For instance, sampling-based methods estimate the marginal effects of a link across multiple sampled subgraphs, while gradient-based methods calculate importance by measuring the sensitivity of task utility to changes in link strength. Finally, communication links often lead to combined causal effects, and their value may not be independent. Group-level perturbation methods evaluate the collective contribution of link clusters, capturing synergistic effects that individual-level estimation may miss. Section E.1 provides further details on potential implementations of this research direction.
@@
One promising direction is to leverage pre-trained language models to generate topologies directly from task descriptions. Given a prompt with task descriptions, agent-role assignments, and interaction patterns, the model could output the micro-level topological structure design by natural language conditioned on intent. This enables rapid adaptation to novel tasks and removes the expensive costs for measuring the effects of each communication link. In addition, Graph generative models, such as graph diffusion [30], conditioned on agent properties, such as roles or functional descriptions, can generate communication topologies that reflect specialization and interdependence. This perspective treats topology construction as a learned mapping from agent features to interaction graphs, enabling compositional and role-aware generation.
4 Challenges and Opportunities in Evaluation
Topological structure learning not only triggers new research directions in methodology and algorithms, but also opens opportunities in evaluation [31, 32]. Current evaluation protocols and systems do not sufficiently satisfy the scenarios in this case. Here, we propose several research directions.
Developing Topology-Aware Benchmarks. There have been plenty of efforts in LLM and agent evaluation [31, 32], primarily focusing on general language understanding, coding, and reasoning tasks. However, more should be done to design topology-aware benchmarks to evaluate MASs in extreme, diverse, and comprehensive scenarios. For instance, how to design a benchmark more suitable for agent selection? How to evaluate the performance if agent numbers are scaling? What data should be collected as the golden standard for validation, given the complex nature of multi-agent systems. Finally, how do we analyze the error propagation in different stages of the topology optimization? We see many open problems and challenges in benchmark design.
Investigating Real Efficacy. While LLMs embedded in multi-agent systems often achieve impressive scores on established benchmarks [33, 34], these metrics may not accurately represent their true capabilities due to potential overfitting and data contamination issues, where models exploit static benchmarks for training rather than demonstrating genuine understanding. It has become critical to determine whether LLMs’ results stem from authentic understanding or mere memorization of training data. Recent research has highlighted that LLM agents can be easily misled by prompts containing incorrect information [35], making them vulnerable to hallucinations and misinformation.
Scaling to Real-world Scenarios. Most benchmark environments test only small-to-medium agent populations, far removed from the scale of real-world systems such as autonomous vehicle fleets or distributed sensor networks. Dynamically Scaling evaluations to hundreds or thousands of agents introduces novel challenges—bandwidth constraints, delayed credit assignment, and intensified non-stationarity-that current setups often ignore. There is a growing need for large-scale, standardized, and dynamic evaluation platforms [36, 37, 38] with reproducible configurations, versioned simulators, and public baselines to facilitate rigorous and comparable research at scale.
Bridging the Simulation-to-Reality Gap. While high-fidelity simulators offer controlled testbeds, they often overlook critical aspects of real-world deployments, such as asynchronous communications, sensor drift, and unmodeled link failures. Evaluation should integrate hardware-in-the-loop testing, field experiments, or shadow deployments to assess performance degradation under operational constraints. Such hybrid evaluation strategies are essential to surface failures that emerge only outside the laboratory and to drive algorithms toward greater robustness in unstructured environments.
Performance Attribution. Understanding why a particular topology performs well is non-trivial. Structural attribution-quantifying the contribution of individual agents or edges—is often confounded by downstream interactions [39]. While counterfactual utility estimation (removing or perturbing agents/edges) is promising, it introduces high computational overhead. Future work should explore group-level attribution techniques that scale with agent count.
5 Discussion: Research in Multiple Systems
While the major content of this paper explores research opportunities for a single system (i.e., a centralized multi-agent architecture), we now discuss emerging research challenges and opportunities in the context of multiple systems—that is, collaborative multi-agent systems distributed across organizations, platforms, or regions. As real-world deployments increasingly involve interconnected services and agent platforms, designing cross-system topological intelligence becomes essential.
Privacy-Preserving Training and Deployment. The growing demand for privacy in LLM systems makes it infeasible to directly share raw data across users, institutions, or geographic boundaries. For example, a personal assistant might interface with separate systems—weather, music, healthcare, finance—each bound by different data governance rules. In such settings, how can we coordinate multiple agent systems while preserving data locality? Future research could explore privacy-aware topological learning, where only encoded summaries or message embeddings are shared. Concepts from federated learning, differential privacy, and secure multi-party computation could be integrated to enable collaboration without information leakage.
Efficient Distributed Optimization. In cross-regional or cross-organizational MAS deployments, agents reside in physically distributed infrastructures. This raises a critical need for latency-aware, topology-optimized communication across system boundaries. How can we dynamically learn which systems should communicate, when, and through which intermediate agents? Opportunities include learning task-conditional, sparse communication bridges, developing hierarchical routing strategies, and combining graph-based scheduling with asynchronous distributed optimization. Moreover, bandwidth-aware structure synthesis and token-efficient inter-system protocols could emerge as a new class of research problems.
Heterogeneous Knowledge Integration. Different systems may specialize in distinct modalities or domains: a hospital system may rely on radiology reports and lab results; a travel planner on temporal and geospatial data; a classroom agent on curriculum structures and student models. These systems exhibit heterogeneous knowledge structures, tool APIs, and communication preferences. The challenge is to design meta-topologies that allow dynamic integration of such capabilities. Future work may explore cross-system role mapping, interface standardization for agent schemas, and structure-aware knowledge alignment techniques. Topology learning in this setting could also facilitate zero-shot system-to-system generalization.
System-Level Alignment and Conflict Resolution. As multiple agent systems interact, they may exhibit divergent objectives, conflicting reasoning paths, or inconsistent internal states. How should topologies be adapted to resolve disagreement across systems? New paradigms in multi-agent negotiation, hierarchical arbitration, and meta-level consensus mechanisms (e.g., judge agents, governance substructures) may need to be integrated. Structure learning in this space will not only optimize for task performance but also for alignment, fairness, and resolution efficiency.
Trust, Robustness, and Verification Across Systems. Multi-system collaborations introduce unique challenges in trust calibration and fault tolerance. If one agent cluster becomes compromised—e.g., via adversarial prompts, faulty tool outputs, or outdated models—how should other systems respond structurally? Research is needed on topological anomaly detection, trust-aware edge pruning, and robust structure adaptation in the presence of partial failures or adversarial agents. Verifiable topologies and explainable structure profiles could become essential for accountability and safety in high-stakes deployments (e.g., finance, healthcare, governance).

### v3

@@ 3.3 Structure Profiling
Empirically, different tasks exhibit distinct coordination patterns. For example, sequential workflows align with chains, hierarchical exploration with trees, centralized tool routing with stars, and high-bandwidth ideation with denser graphs Zhang et al. (2024e); Yao et al. (2023); Besta et al. (2024). The macro-level topological structure in MASs directly determines the economic burden of collaboration. Imagining all agents in MASs adopt a fully connected structure, each round of interaction generates a large volume of redundant message exchanges, which not only increases token consumption but also leads to significant latency Wu et al. (2023). This issue becomes even more pronounced as the system scales, where MAS faces the risk of exponential growth in token usage. For example, experiments demonstrate that unoptimized multi-round fully connected MAS can consume tens of times more tokens than single-agent or chain-structured systems Wang et al. (2025). Committing early to a macro-level archetype gives the system a strong inductive bias before any micro-level wiring, improving efficiency, robustness, and interpretability across tasks. We frame this explicitly as selecting among a discrete set of topological archetypes conditioned on the task, rather than directly optimizing the full edge set. Archetype selection injects the right inductive bias upfront (e.g., depth for decomposition, breadth for exploration), which stabilizes subsequent micro-search by bounding the feasible graph family and reducing non-stationarity in edge learning. Therefore, given a task $x\in T$, the goal is to select a single macro topology $M$ from a pre-defined structure pool $\mathcal{M}$ (e.g., chain, tree, star, ring, fully-connected) that best trades off expected task reward and communication budget Wang et al. (2025); Besta et al. (2024). We adopt a cost-aware criterion to pick the macro topology:
$\vskip-5.0ptM^{*}\;=\;\arg\max_{M\in\mathcal{M}}\;\;\mathcal{R}(x,M)\;-\;\lambda\,C(M),\vskip-5.0pt$
where $\mathcal{R}(x,M)$ estimates expected task reward under topology $M$ and $C(M)$ aggregates communication-related costs (e.g., token, latency), with $\lambda$ balancing quality and economy. The pool should be a small, human-interpretable set and can be expanded with DAG variants or layered stars when justified by use cases Zhang et al. (2024e).
For structure profiling, a compact classifier $f_{\theta}(\phi(T))\to\mathcal{M}$ on (task, structure) pairs curated from prior runs, simulations, or modest expert annotation should be a feasible way, where task signature $\phi(T)$ is a descriptor derived from the prompt, tools, constraints, or historical traces (e.g., need for depth or breadth, central routing or peer debate). Using cross-entropy with cost-aware reweighting to reflect $C(M)$ (e.g., penalizing dense graphs) should also be considered Wu et al. (2023); Chen et al. (2023). The output will be $M^{*}$ with possible macro hyper-parameters, such as max degree, allowed diameter, or layer count).
@@
How do we learn mappings from task signatures to a small set $\mathcal{M}$ that transfer across domains, tools, and models? We may need strategies that leverage weak supervision, priors, and continual updates without drifting. (P1, P2)
Open Problem 4.
Different macro-level topological structures in MASs can lead to drastically different communication costs. How can we control the resource cost while maintaining the task performance? (P1)
@@
Design few-shot structure classifiers with uncertainty. The compact classifiers can be trained on modest (task, archetype) corpora, with uncertainty gating to trigger budgeted probes only when needed; continual learning from deployment logs.
The robustness of MASs critically depends on their topological structure, which shapes the system’s tolerance to failures and adversarial influences Wu et al. (2023); Yao et al. (2023). For example, tree structures may provide cross-verification, allowing the system to recover when individual agents fail or inject misinformation. In contrast, chain structures are fragile, as the failure of a single agent can break the reasoning flow, and adversarial perturbations introduced early can propagate unchecked through subsequent steps Du et al. (2023); Hong et al. (2023). These observations underscore that robustness is inherently a macro-level topological property, aligning with our position P3.
Open Problem 5.
What redundancy is minimally sufficient to resist agent failures and adversaries without prohibitive overhead, and how do macro choices affect fair access to central nodes (e.g., star) and bias amplification? (P3, P4)
@@
Benchmarks will be valuable and should annotate macro labels as well as report metrics by archetype class so we can attribute gains to topology rather than model scale or prompt tricks.
The structure profiling stabilizes subsequent micro-search, and enables principled cost–quality trade studies. Next, Section 3.4 will instantiate edges and communication policies strictly within those bounds.
@@ 3.4 Topological Structure Synthesis
While agent selection and structure profiling identify the macro-level pattern of interactions, synthesizing a micro-level communication topology is a pressing problem to further enhance the communication efficiency and task performance. Specifically, agent selection determines who participates in the collaboration, and structure profiling identifies the macro-level pattern of interaction. Effectively executing the task requires synthesizing a fine-grained communication topology that governs the actual flow of information between agents. This micro-level structure plays a critical role in determining not only which agents should communicate but also the edge directions, weights, and potential hierarchy (e.g., tree depth or breadth). These decisions directly influence how efficiently information propagates throughout the system, ultimately shaping task performance (Wang et al., 2025), communication cost (Wang et al., 2025; Zhang et al., 2024a), and the robustness of the multi-agent collaboration (Yu et al., 2024).
@@
Fully connecting every pair of agents leads to redundancy and high communication costs: How do we prune unnecessary links while preserving task performance? (P1)
Recent efforts have explored topological structure optimization in MASs through topology pruning-based methods, which aim to reduce redundancy and improve efficiency by selectively removing communication links from the topological structure. AgentPrune (Zhang et al., 2024a) learns a sparse spatial-temporal communication graph by optimizing differentiable edge masks with policy gradients and performs one-shot pruning to remove redundant message-passing edges while preserving overall task performance. Similarly, G-Designer (Zhang et al., 2024b) adopts a reinforcement learning-based framework to learn task-specific sparse communication graphs by selecting critical communication links. AgentDropout (Wang et al., 2025), in contrast, adaptively drops low-utility intra- and inter-round edges to compress communication graphs and reduce token usage during inference. GPTSwarm (Zhuge et al., 2024) regards agents as nodes in a computational graph and optimizes inter-agent communication links, enabling end-to-end learning of sparse collaboration topologies tailored to task requirements. However, most of them suffer from limited scalability and high computational cost, leaving an open problem that urgently needs to be addressed in future research, which directly verifies our position P1.
Open Problem 7.
Different agent roles should occupy specific positions within the topological structure: How do we design the micro-structure while considering the specific role of each agent to maximize performance, make it more robust, while maintaining the fairness? (P2, P3, P4)
While many existing approaches focus on the communication connectivity among agents Zhang et al. (2024a), they often ignore the problem of the different, specialized roles of agents that should occupy distinct and meaningful positions within the topology. For example, a manager should be in the central position for overall coordination, while a verifier may serve best at the end of the pipeline. Therefore, this open problem implies the position statement of P2.
In addition, several existing studies have begun to explore the robustness of different topological structures in MASs Yu et al. (2024). For example, in some cases, a manager placed at the center can maintain coordination even if peripheral agents fail, while a verifier positioned at the end of the pipeline ensures that adversarial perturbations do not propagate unchecked. Thereby, once the specific role of each agent is aligned with the task’s requirements, robustness emerges as a central challenge, which is exactly the concern articulated in our position P3.
Lastly, fairness is also deeply intertwined with the placement of agent roles. If certain roles (e.g., managers or planners) are always granted central or high-degree positions, while others (e.g., verifiers or critics) are pushed to the periphery, the system risks privileging dominant voices and marginalizing supporting ones. A fair topology should instead ensure that diverse roles and personas have equitable opportunities to contribute, preventing structural biases from distorting decision-making. This directly connects to our position P4.
@@
A central challenge in optimizing micro-level topologies in MASs is to assess each communication link’s contribution to task performance. Beyond coarse assumptions (e.g., fully connected), fine-grained evaluation enables pruning strategies that reduce overhead while preserving accuracy. Counterfactual reasoning estimates link utility from performance drops upon removal, while scalable approximations such as sampling or gradient sensitivity offer efficiency. Since links interact, group-level perturbations capture collective contributions that individual assessments may miss. Section˜E.1 provides further details on potential implementations of this research direction.
@@
Pre-trained language models can directly generate micro-level topologies from task descriptions, avoiding costly link-level evaluation. Likewise, graph generative models conditioned on agent roles can produce role-aware interaction structures that capture specialization and interdependence.
4 Challenges and Discussions
The positions and research directions articulated above are both compelling and necessary. Yet, despite their clarity, most MASs still fall short of realizing them in practice. We believe the following constitute the principal barriers:
Developing Topology-Aware Benchmarks. Current benchmarks focus on general reasoning or coding tasks and fail to capture topology-related dynamics such as agent scaling, communication redundancy, and error propagation. Future work should design topology-aware benchmarks and metrics that reflect coordination efficiency and structure sensitivity.
Real-World Robustness. While many MASs achieve high benchmark scores, they often rely on static or contaminated data. Evaluating genuine reasoning capability under noisy, adversarial, or evolving environments remains difficult.
Scaling and Reality Gap. Most MASs are tested on small populations within simulators. Scaling to hundreds of agents or bridging the simulation-to-reality gap introduces latency, bandwidth, and stability issues that current setups overlook.
Multi-System Coordination. In real deployments, MASs span multiple organizations and data boundaries. This raises challenges in privacy-preserving topology learning, distributed optimization, and cross-system alignment.
Trust and Accountability. As MAS ecosystems grow, ensuring verifiable, trustworthy, and failure-resilient topologies becomes essential. Research on trust-aware pruning, structural verification, and explainable topology design is urgently needed.
Expensive Optimization. Most topology optimization methods for MASs are computationally expensive, such as relying on counterfactual approaches. Their cost grows rapidly with the number of agents and links, making it unrealistic.
@@ 5 Conclusion
6 Limitations
While this paper advocates for topological structure learning as a critical research priority for LLM-based MASs, our discussion remains primarily conceptual and forward-looking. In addition, the proposed three-stage framework is summarized from patterns observed across most existing works, serving as a generalized abstraction rather than an exhaustive taxonomy. However, certain specialized or hybrid optimization methods may not fit neatly into this framework, which could limit the completeness of our argumentation for the stated positions. Last but not least, our analysis primarily centers on LLM-based textual agents. Extending the framework to multimodal or embodied agents (e.g., vision-language or robotics systems) may introduce additional complexities not addressed in this paper.
7 Ethical Considerations
Optimizing the topological structure of MASs may lead to new ethical risks beyond efficiency and performance. Structural imbalances could inadvertently privilege certain agents, underscoring the importance of fairness-aware topology learning. Besides, cross-system communication also requires careful treatment of privacy and governance boundaries.